OpenStudy (anonymous):

A box of colored thumb tacks contains 80 tacks with an equal assortment of red green blue yellow tacks. If you close your eyes and select a single thumb tack, what is the probability that it will NOT be red? Enter answer in simplest form

OpenStudy (anonymous):

it is 1 - P(R), where P(R) is the probability that the single thumb tack you picked was red P(R) = (80 * 0.25)/80 = 0.25 = 1/4 so 1- 0.25 = 0.75

OpenStudy (anonymous):

wats the answer in simplest form again??? is it 1/4... the answer hasto be a fraction in simplest form

OpenStudy (anonymous):

nevermind its 3/4